Good evening, I am taking an Algebra class for the first time after 50 years, and can not remember how to find the point-slope form and the slope-intersect form of a line for the following problem. M=-1/5,P(10,4)
Any assistance would be greatly appreciated.
Sarita

Found 2 solutions by FrankM, ikleyn:
Answer by FrankM(1040)

Point slope is
Y-Y1= m (X-X1) where m is the slope, and X1,Y1 is any point on the line.
Y-4=-1/5 (X-10) is the point slope form.
Distribute -1/5
Y-4= -1/5X+2
Y=-1/5X+6 slope-intercept.

(And thank you, for inviting me to post my 1000th answer on this site!)

Answer by ikleyn(52787)

.
There is a group of lessons in this site that explain this situation and all closely related other cases

    - Find the slope of a straight line in a coordinate plane passing through two given points
    - Equation for a straight line having a given slope and passing through a given point
    - Solving problems related to the slope of a straight line
    - Equation for a straight line in a coordinate plane passing through two given points
    - Equation for a straight line parallel to a given line and passing through a given point
    - Equation for a straight line perpendicular to a given line and passing through a given point

You may learn everything from these lessons once and for all !


Or use it as your "handbook" and a reference !